# Revolutionizing Future Connectivity: A Contemporary Survey on AI-empowered Satellite-based Non-Terrestrial Networks in 6G

### v3

@@ II-A2 Air-borne platforms
High Altitude Platform Systems (HAPS) refer to air-borne platforms that can be used for wireless communication. Airships, balloons, and airplanes are the most prominent types of air-borne platforms in NTN. They are viewed as air-borne counterparts of terrestrial base stations serving as High-altitude International Mobile Base Stations (HIBS) [21]. They usually operate at the stratosphere region with an altitude of around 20 km and a beam footprint size with a diameter of several km. Despite it having a lot smaller propagation delay compared to space-borne platforms, it has some additional challenges related to stabilization on air and refueling.
@@ II-B Role of NTN in 6G
NTNs are anticipated to be a major component of 6G communication systems, providing a wide array of vertical services, such as transport, health, energy, automotive, public safety, and many more. The International Telecommunication Union (ITU) has identified three major categories of applications for 5G that are based on network performance and user Quality of Experience (QoE): (1) eMBB: extremely high bandwidth with moderate latency requirements, for example, multimedia applications; (2) mMTC: low power and bandwidth and no strict delay requirements, for example, IoT; and (3) URLLC: low latency and high-reliability requirements, for example, remote medical surgery. However, future applications such as AR, VR, Tactile Internet, HTC, intelligent transport and automation, multi-sense communication, global ubiquitous connectivity, etc. require extremely high throughput, low latency, high reliability, and ubiquity at the same time which cannot be met with current 5G standards [4]. Based on the characteristics, these new applications are classified into three more new groups,
@@ II-B2 Ultra-reliable low Latency Broadband Communication (ULBC)
The principal strength of NTNs lies in their extreme coverage. As discussed in Hexa-X project [22], a flagship for B5G/6G vision and intelligent fabric of technology enablers connecting human, physical, and digital worlds, the vision of enabling 6G networks towards provisioning service everywhere and always through NTN is presented. Due to its extreme coverage, satellites can reach underserved or unserved areas such as islands, remote locations, ships, airplanes, etc. where terrestrial communication is either difficult or impossible to some extent. In times of natural disaster, terrestrial links can be unavailable, in which case users can benefit from the reliable backup of non-terrestrial links. This ensures resilient and robust communication with global connectivity which is considered to be one of the main features of future 6G networks. With the advancements in antenna techniques and miniaturization, high throughput satellites are also deployed in low earth orbits. Consequently, current 5G use cases such as mMTC and eMBB as well as future 6G use cases such as uMBB can be the most important use cases for NTNs. Furthermore, the considerably low latency for LEO satellite systems makes the satellite useful even for low-latency applications. However, 5G URLLC or 6G new use cases with extremely low latency may not be directly applicable for NTN use cases. Nevertheless, NTNs can still be beneficial for these use cases in conjunction with terrestrial networks to improve network efficiencies and reliability. Combining all these, satellites are expected to be one of the major driving forces toward revolutionizing the future 6G applications extensively.
@@ II-C General Architecture
Satellites can employ a transparent payload configuration, acting as a relay that performs RF filtering, frequency conversion, and amplification to facilitate communication between UEs and ground stations. Alternatively, they can utilize a regenerative payload configuration, which involves payload processing after modulation and coding, and act as base stations with additional onboard processing capabilities. Besides, the satellites can provide backhaul support for the core networks of terrestrial networks. The general architecture for a satellite-based NTN for the above-mentioned different configurations as per release 16 and 17 is discussed below [20, 23]:
@@
Satellite: Satellite is the key component of this architecture. It carries the payload between the UE and the ground station as shown in Figure 3. In the case of a transparent payload, it works as a simple relay that transmits the payload after RF filtering, frequency conversion, and amplification to the ground station (or UE). Conversely, in the case of a regenerative payload, it processes the payload after modulation and coding on top of these actions, so it works like a BS that needs onboard processing capabilities. Also as per [23], satellites can provide backhaul by providing a connection between ground BS and the core network as illustrated in Figure 3.
@@ II-D1 Target Frequency Band
The allowable frequency range of operation is 0.5-100 GHz [24]. Traditionally, six major frequency bands within this range are used in satellite communications, which are listed below:
@@
However, as per 3GPP, currently, two frequency bands (S and Ka-band) are targeted in particular for integrated TNTN environments considering performance and regulatory concerns [20]. The two target frequency bands are:
@@ II-D2 Propagation Delay
Propagation delay is the time duration taken for a signal to reach its destination. For communication signals, we can calculate the propagation delay for a signal by using the equation: where is the distance between the source and destination and m/s is the speed of light. Considering the speed of light as a constant, we observe the propagation delay for a signal is proportional to the propagation distances. Satellites are located very far from the surface of the earth as discussed in Section II-A. Consequently, the propagation delay is going to be extremely large for NTNs. The GEO, MEO, and LEO satellites can have a one-way propagation delay of about 270, 94, and 20 ms respectively as shown in Table II. These values are much larger, especially for GEO and MEO cases compared to conventional terrestrial networks, which generally have a very negligible propagation delay of around a few [20]. This extended propagation delay has an effect on different network procedures and performances for communication systems.
@@ II-D3 Propagation Loss
The propagation loss, or path loss, refers to the reduction in power density that an electromagnetic signal experiences as it travels through space. The most significant component of this path loss is the free space path loss, which is proportional to the distance between the source and destination and the frequency of the signal [25]. For NTNs, this free space path loss is much higher (around 60-120 dB) than it is for terrestrial networks, due to the greater distances between satellites and the use of higher carrier frequencies. In fact, the Ka-band is not suitable for GEO satellites, as it does not meet the minimum link budget for them. The basic path loss component also includes shadow fading [26], as with traditional terrestrial networks.
In addition to that, there is attenuation due to atmospheric gases that depends on frequency, elevation angle, altitude above sea level, and water vapor density[27]. Another important component is attenuation due to rain and fog, which is typically significant for frequencies above GHz [28]. Additionally, scintillation corresponds to rapid fluctuation in amplitude and phase of the propagating signal in the ionosphere (for Sub- GHz) and troposphere (for above GHz) [29]. Depending on different scenarios, either flat fading based on ITU two-state model [30] or fast fading [20] can be considered. The average propagation loss for different types of satellites in different frequency bands is illustrated in Figure 4. This high path loss necessitates the need for efficient power allocation strategies in NTN.
@@ II-D5 Coverage Area
One of the most important features of the satellites is the large beam footprint associated with them due to their long distances from the earth’s surface. This enables the network coverage of very large areas compared to the coverage area of terrestrial counterparts. It provides us with ubiquitous network coverage including remote, even isolated areas. However, this also creates the necessity for modifications in existing timing and synchronization procedures for conventional terrestrial networks. The cell area is much larger, so the UEs situated at the farthest side of the cells experience a larger delay compared to the UEs situated closer to the satellites [20]. So the timestamps for different network procedures need to be modified according to the distances of the users as we will see in the next subsection.
@@ II-E1 Channel Estimation
In wireless communications, Channel State Information (CSI) refers to the information which represents the state of a communication channel between the transmitter(s) and the receiver(s); the process of obtaining this information is known as channel estimation. By having access to CSI, it is possible to adjust transmissions to the current channel conditions, which is essential for achieving reliable communication with high data rates in multi-antenna systems with effective channel resources and interference management. There are numerous advanced approaches, such as Maximum Likelihood estimation and Minimum Mean Square Error (MMSE) estimation, for effective channel estimation in traditional terrestrial cellular networks. Nevertheless, these methods are not well-suited for NTN, particularly for LEO satellites due to the inherent time-variant nature of the satellite communication channels. LEO satellites usually move from horizon to horizon in approximately 5-10 minutes, so a UE remains within the coverage of a specific LEO satellite for a very short time period. Furthermore, the propagation delay for satellite networks, especially in the case of GEO satellites, is considerably larger (250 ms RTT) in comparison to general terrestrial networks. Therefore, the CSI estimated by the LEO satellites can be outdated [31]. Because of these reasons, the CSI estimation in NTN necessitates new efficient techniques in addition to the traditional terrestrial estimation methods.
@@ II-E2 Mobility Management
Since an NGEO satellite operates at a lower altitude, the coverage area of each NGEO satellite is smaller than that of a GEO satellite. Typically around 5-20, NGEO satellites form complex mega-constellations to sustain global coverage across the earth. The NGEO satellite needs to move at a much higher speed than the earth’s rotational speed (can be up to around 7.8 km/s) to get the necessary centripetal force to move around the earth at that low altitude. As a result, these satellites typically orbit around the earth pretty fast (usually within around 2-10 hours) as discussed in Section II-A. This quick orbital motion poses a great challenge for integrating NGEO satellites into traditional wireless communication systems. Due to the smaller orbital period, any specific terrestrial UE can be only visible to an NGEO satellite for a very short span of time, typically several minutes. So the UE needs to undergo multiple handovers within a short span of time interval regardless of its mobility [14]. If the satellite covers an area using multiple spot beams, the scenario is worse because the spot beam is much smaller compared to a total coverage area of an NGEO satellite. So the UEs need to through multiple (beam) handovers within a few minutes, even when they are stationary, for seamless continuation of data sessions. This frequent handover phenomenon in NGEO satellite networks creates a lot of overhead in communication channels, leading to an overall degradation in network performance.
@@ II-E3 High Doppler Shift
Doppler shift is the shift in the signal frequency due to the motion of the transceivers. In the case of NGEO satellites, satellites are moving at a very high speed under a specific constellation. Due to this relative motion between UEs and satellites, Doppler shift happens in the original signal frequency. Due to frequency offsets, UEs tune to different carrier frequencies than the original carrier frequencies. So the frequency synchronization is lost, and the UEs may interfere with the other users. This is known as Inter-Carrier Interference (ICI) between multiple UEs. Generally, even for the high mobility scenarios in terrestrial networks, the frequency shift is pretty negligible, and so is the Doppler shift. However, the frequency offset is pretty significant in NTN due to the much higher speed of the NGEO satellites. The Doppler shift value mainly depends on the carrier frequency and height of the satellites. For NGEO satellites operating at Ka-Band, the Doppler shift can go from 225 kHz to even 720 kHz [20] depending on the heights. This can cause significant ICI among NTN users which requires efficient strategies for compensation of the Doppler effect.
@@ II-E6 Effect on Network Procedures
Timing advances ensure synchronous uplink transmissions for all UEs. The UEs can be located at different distances from the gNB, so there is a differential propagation delay between different UEs. If the uplink reception is not synchronized, the gNB needs to make sure the allocation of resource blocks to a specific UE does not include the resource blocks already in use by other UEs, which is inefficient in terms of resource allocation. Due to the long propagation delay, the TA is much larger than the transmission time slots in NTN compared to NR. Also due to the mobility of LEO satellites, the delay is time-varying and TA needs dynamic updates for proper uplink alignments. The other processes affected by the long propagation delay are Random Access, Hybrid Automatic Repeat Requests (HARQ) procedures, etc [20]. These procedures need to be modified properly to compensate for the long propagation delay.
@@ III AI and Its Relevance to NTN Challenges
AI refers to the simulation of human intelligence processes (e.g. visual reception, speech recognition, computer vision, etc.) by machines, especially computer systems. This human-level cognitive ability is achieved through either some predefined algorithms or learning from data-based approaches [15]. Many practical systems are very diverse and complex. The rule-based approaches are not very feasible for these systems because of an enormous number of scenario possibilities. As a result, the learning-based approaches show a lot more promise compared to predefined approaches in these types of real systems. As our focus for this paper is mostly on NTN which has an extremely complex and time-variant topology, we focus on learning-based approaches when we consider AI. In this section, we give an overview of these approaches to get an intuition of how these approaches can be useful in solving NTN issues discussed in the next section.
@@ III-A Machine Learning (ML)
Machine Learning (ML) is a special subset of AI approaches where machines learn algorithms to perform a task by generalizing from past experiences or historical data without being explicitly programmed for it [32]. The performance of human intelligence processes can be improved with each iteration evolving through new feature extractions in ML approaches. Generally speaking, each ML approach has three distinctive features, namely task, performance measure, and experience [33]. A machine is first assigned to learn to perform a specific task. It starts with a model with an initial set of random parameters. Then at every iteration, the model is recalculated based on some performance measures, essentially representing the learning process. Thus utilizing experiences, it can learn how to perform the task properly which is the main goal of ML approaches.
A generic ML model works in three phases utilizing various components [16] - Pre-Training Phase, Training Phase, and Testing Phase. We discuss the fundamental components of these three phases as shown in Figure 5 below:
@@ III-C Deep Learning (DL)
In complex real-world problems, feature extraction can turn out to be extremely challenging using generic ML models. There may be hundreds of parameters that need to be learned and the outputs may not be linearly correlated to the inputs. So general ML models may not provide satisfactory performance in learning these problems. To facilitate mapping outputs to inputs, Neural Networks (NNs) [34] are widely used in ML frameworks. With the availability of a large amount of data, NNs have emerged as a key technology to be used in ML in the recent past. The learning process can be largely benefited from the introduction of NNs to deal with complicated large-scale problems. This learning process involving NNs to estimate the models is known as Deep Learning (DL) [17] which is a special important subset of ML.
NNs are inspired by the biological neural networks in the brain, more specifically the nervous system. To mimic the operation of the brain, the NNs are composed of multiple layers where each layer consists of multiple neurons followed by an activation function. Generally, the neurons in one layer are connected to the neurons in the adjacent layers. The connecting edges have weights that represent the relationship between the neurons. Each layer output can be viewed as some intermediate decisions which eventually result in the final output values. The weights are generally trained through a number of iterations using backpropagation algorithms [35]. Generally, the cost function associated with the model to calculate the difference between the predicted and actual outputs is not very simple, so we use different numerical methods like Gradient Descent [36], Stochastic Gradient Descent [37], Mini-Batch Stochastic Gradient Descent [38], Newton’s method [39] etc. and so on to estimate the gradients of the cost function with respect to corresponding weights. At each iteration, the weights are updated by an amount based on these calculated gradients and a predefined learning rate. As we move towards the gradient descent direction, it helps us to reduce the cost at every iteration. In this manner, we can map the inputs to outputs through NNs.

### v4

@@ II-A2 Air-borne platforms
High Altitude Platform Systems (HAPS) refer to air-borne platforms that can be used for wireless communication. Airships, balloons, and airplanes are the most prominent types of air-borne platforms in NTN. They are viewed as air-borne counterparts of terrestrial base stations serving as High-altitude International Mobile Base Stations (HIBS) [haps]. They usually operate at the stratosphere region with an altitude of around 20 km and a beam footprint size with a diameter of several km. Despite it having a lot smaller propagation delay compared to space-borne platforms, it has some additional challenges related to stabilization on air and refueling.
@@ II-B Role of NTN in 6G
NTNs are anticipated to be a major component of 6G communication systems, providing a wide array of vertical services, such as transport, health, energy, automotive, public safety, and many more. The International Telecommunication Union (ITU) has identified three major categories of applications for 5G that are based on network performance and user Quality of Experience (QoE): (1) eMBB: extremely high bandwidth with moderate latency requirements, for example, multimedia applications; (2) mMTC: low power and bandwidth and no strict delay requirements, for example, IoT; and (3) URLLC: low latency and high-reliability requirements, for example, remote medical surgery. However, future applications such as AR, VR, Tactile Internet, HTC, intelligent transport and automation, multi-sense communication, global ubiquitous connectivity, etc. require extremely high throughput, low latency, high reliability, and ubiquity at the same time which cannot be met with current 5G standards [6g-1]. Based on the characteristics, these new applications are classified into three more new groups,
@@ II-B2 Ultra-reliable low Latency Broadband Communication (ULBC)
The principal strength of NTNs lies in their extreme coverage. As discussed in Hexa-X project [hexax1], a flagship for B5G/6G vision and intelligent fabric of technology enablers connecting human, physical, and digital worlds, the vision of enabling 6G networks towards provisioning service everywhere and always through NTN is presented. Due to its extreme coverage, satellites can reach underserved or unserved areas such as islands, remote locations, ships, airplanes, etc. where terrestrial communication is either difficult or impossible to some extent. In times of natural disaster, terrestrial links can be unavailable, in which case users can benefit from the reliable backup of non-terrestrial links. This ensures resilient and robust communication with global connectivity which is considered to be one of the main features of future 6G networks. With the advancements in antenna techniques and miniaturization, high throughput satellites are also deployed in low earth orbits. Consequently, current 5G use cases such as mMTC and eMBB as well as future 6G use cases such as uMBB can be the most important use cases for NTNs. Furthermore, the considerably low latency for LEO satellite systems makes the satellite useful even for low-latency applications. However, 5G URLLC or 6G new use cases with extremely low latency may not be directly applicable for NTN use cases. Nevertheless, NTNs can still be beneficial for these use cases in conjunction with terrestrial networks to improve network efficiencies and reliability. Combining all these, satellites are expected to be one of the major driving forces toward revolutionizing the future 6G applications extensively.
@@ II-C General Architecture
Satellites can employ a transparent payload configuration, acting as a relay that performs RF filtering, frequency conversion, and amplification to facilitate communication between UEs and ground stations. Alternatively, they can utilize a regenerative payload configuration, which involves payload processing after modulation and coding, and act as base stations with additional onboard processing capabilities. Besides, the satellites can provide backhaul support for the core networks of terrestrial networks. The general architecture for a satellite-based NTN for the above-mentioned different configurations as per release 16 and 17 is discussed below [3gpp1, 3gpp3]:
@@
Satellite: Satellite is the key component of this architecture. It carries the payload between the UE and the ground station as shown in Figure 3. In the case of a transparent payload, it works as a simple relay that transmits the payload after RF filtering, frequency conversion, and amplification to the ground station (or UE). Conversely, in the case of a regenerative payload, it processes the payload after modulation and coding on top of these actions, so it works like a BS that needs onboard processing capabilities. Also as per [3gpp3], satellites can provide backhaul by providing a connection between ground BS and the core network as illustrated in Figure 3.
@@ II-D1 Target Frequency Band
The allowable frequency range of operation is 0.5-100 GHz [spec-shar]. Traditionally, six major frequency bands within this range are used in satellite communications, which are listed below:
@@
However, as per 3GPP, currently, two frequency bands (S and Ka-band) are targeted in particular for integrated TNTN environments considering performance and regulatory concerns [3gpp1]. The two target frequency bands are:
@@ II-D2 Propagation Delay
Propagation delay is the time duration taken for a signal to reach its destination. For communication signals, we can calculate the propagation delay for a signal by using the equation: where is the distance between the source and destination and m/s is the speed of light. Considering the speed of light as a constant, we observe the propagation delay for a signal is proportional to the propagation distances. Satellites are located very far from the surface of the earth as discussed in Section II-A. Consequently, the propagation delay is going to be extremely large for NTNs. The GEO, MEO, and LEO satellites can have a one-way propagation delay of about 270, 94, and 20 ms respectively as shown in Table II. These values are much larger, especially for GEO and MEO cases compared to conventional terrestrial networks, which generally have a very negligible propagation delay of around a few [3gpp1]. This extended propagation delay has an effect on different network procedures and performances for communication systems.
@@ II-D3 Propagation Loss
The propagation loss, or path loss, refers to the reduction in power density that an electromagnetic signal experiences as it travels through space. The most significant component of this path loss is the free space path loss, which is proportional to the distance between the source and destination and the frequency of the signal [ntn-fspl]. For NTNs, this free space path loss is much higher (around 60-120 dB) than it is for terrestrial networks, due to the greater distances between satellites and the use of higher carrier frequencies. In fact, the Ka-band is not suitable for GEO satellites, as it does not meet the minimum link budget for them. The basic path loss component also includes shadow fading [ntn-sf], as with traditional terrestrial networks.
In addition to that, there is attenuation due to atmospheric gases that depends on frequency, elevation angle, altitude above sea level, and water vapor density[ntn-atm]. Another important component is attenuation due to rain and fog, which is typically significant for frequencies above GHz [ntn-rf]. Additionally, scintillation corresponds to rapid fluctuation in amplitude and phase of the propagating signal in the ionosphere (for Sub- GHz) and troposphere (for above GHz) [ntn-sci]. Depending on different scenarios, either flat fading based on ITU two-state model [ntn-ts] or fast fading [3gpp1] can be considered. The average propagation loss for different types of satellites in different frequency bands is illustrated in Figure 4. This high path loss necessitates the need for efficient power allocation strategies in NTN.
@@ II-D5 Coverage Area
One of the most important features of the satellites is the large beam footprint associated with them due to their long distances from the earth’s surface. This enables the network coverage of very large areas compared to the coverage area of terrestrial counterparts. It provides us with ubiquitous network coverage including remote, even isolated areas. However, this also creates the necessity for modifications in existing timing and synchronization procedures for conventional terrestrial networks. The cell area is much larger, so the UEs situated at the farthest side of the cells experience a larger delay compared to the UEs situated closer to the satellites [3gpp1]. So the timestamps for different network procedures need to be modified according to the distances of the users as we will see in the next subsection.
@@ II-E1 Channel Estimation
In wireless communications, Channel State Information (CSI) refers to the information which represents the state of a communication channel between the transmitter(s) and the receiver(s); the process of obtaining this information is known as channel estimation. By having access to CSI, it is possible to adjust transmissions to the current channel conditions, which is essential for achieving reliable communication with high data rates in multi-antenna systems with effective channel resources and interference management. There are numerous advanced approaches, such as Maximum Likelihood estimation and Minimum Mean Square Error (MMSE) estimation, for effective channel estimation in traditional terrestrial cellular networks. Nevertheless, these methods are not well-suited for NTN, particularly for LEO satellites due to the inherent time-variant nature of the satellite communication channels. LEO satellites usually move from horizon to horizon in approximately 5-10 minutes, so a UE remains within the coverage of a specific LEO satellite for a very short time period. Furthermore, the propagation delay for satellite networks, especially in the case of GEO satellites, is considerably larger (250 ms RTT) in comparison to general terrestrial networks. Therefore, the CSI estimated by the LEO satellites can be outdated [3gpp2]. Because of these reasons, the CSI estimation in NTN necessitates new efficient techniques in addition to the traditional terrestrial estimation methods.
@@ II-E2 Mobility Management
Since an NGEO satellite operates at a lower altitude, the coverage area of each NGEO satellite is smaller than that of a GEO satellite. Typically around 5-20, NGEO satellites form complex mega-constellations to sustain global coverage across the earth. The NGEO satellite needs to move at a much higher speed than the earth’s rotational speed (can be up to around 7.8 km/s) to get the necessary centripetal force to move around the earth at that low altitude. As a result, these satellites typically orbit around the earth pretty fast (usually within around 2-10 hours) as discussed in Section II-A. This quick orbital motion poses a great challenge for integrating NGEO satellites into traditional wireless communication systems. Due to the smaller orbital period, any specific terrestrial UE can be only visible to an NGEO satellite for a very short span of time, typically several minutes. So the UE needs to undergo multiple handovers within a short span of time interval regardless of its mobility [NTN_2]. If the satellite covers an area using multiple spot beams, the scenario is worse because the spot beam is much smaller compared to a total coverage area of an NGEO satellite. So the UEs need to through multiple (beam) handovers within a few minutes, even when they are stationary, for seamless continuation of data sessions. This frequent handover phenomenon in NGEO satellite networks creates a lot of overhead in communication channels, leading to an overall degradation in network performance.
@@ II-E3 High Doppler Shift
Doppler shift is the shift in the signal frequency due to the motion of the transceivers. In the case of NGEO satellites, satellites are moving at a very high speed under a specific constellation. Due to this relative motion between UEs and satellites, Doppler shift happens in the original signal frequency. Due to frequency offsets, UEs tune to different carrier frequencies than the original carrier frequencies. So the frequency synchronization is lost, and the UEs may interfere with the other users. This is known as Inter-Carrier Interference (ICI) between multiple UEs. Generally, even for the high mobility scenarios in terrestrial networks, the frequency shift is pretty negligible, and so is the Doppler shift. However, the frequency offset is pretty significant in NTN due to the much higher speed of the NGEO satellites. The Doppler shift value mainly depends on the carrier frequency and height of the satellites. For NGEO satellites operating at Ka-Band, the Doppler shift can go from 225 kHz to even 720 kHz [3gpp1] depending on the heights. This can cause significant ICI among NTN users which requires efficient strategies for compensation of the Doppler effect.
@@ II-E6 Effect on Network Procedures
Timing advances ensure synchronous uplink transmissions for all UEs. The UEs can be located at different distances from the gNB, so there is a differential propagation delay between different UEs. If the uplink reception is not synchronized, the gNB needs to make sure the allocation of resource blocks to a specific UE does not include the resource blocks already in use by other UEs, which is inefficient in terms of resource allocation. Due to the long propagation delay, the TA is much larger than the transmission time slots in NTN compared to NR. Also due to the mobility of LEO satellites, the delay is time-varying and TA needs dynamic updates for proper uplink alignments. The other processes affected by the long propagation delay are Random Access, Hybrid Automatic Repeat Requests (HARQ) procedures, etc [3gpp1]. These procedures need to be modified properly to compensate for the long propagation delay.
@@ III AI and Its Relevance to NTN Challenges
AI refers to the simulation of human intelligence processes (e.g. visual reception, speech recognition, computer vision, etc.) by machines, especially computer systems. This human-level cognitive ability is achieved through either some predefined algorithms or learning from data-based approaches [ai]. Many practical systems are very diverse and complex. The rule-based approaches are not very feasible for these systems because of an enormous number of scenario possibilities. As a result, the learning-based approaches show a lot more promise compared to predefined approaches in these types of real systems. As our focus for this paper is mostly on NTN which has an extremely complex and time-variant topology, we focus on learning-based approaches when we consider AI. In this section, we give an overview of these approaches to get an intuition of how these approaches can be useful in solving NTN issues discussed in the next section.
@@ III-A Machine Learning (ML)
Machine Learning (ML) is a special subset of AI approaches where machines learn algorithms to perform a task by generalizing from past experiences or historical data without being explicitly programmed for it [ai-sam]. The performance of human intelligence processes can be improved with each iteration evolving through new feature extractions in ML approaches. Generally speaking, each ML approach has three distinctive features, namely task, performance measure, and experience [ai-mit]. A machine is first assigned to learn to perform a specific task. It starts with a model with an initial set of random parameters. Then at every iteration, the model is recalculated based on some performance measures, essentially representing the learning process. Thus utilizing experiences, it can learn how to perform the task properly which is the main goal of ML approaches.
A generic ML model works in three phases utilizing various components [ai-ml] - Pre-Training Phase, Training Phase, and Testing Phase. We discuss the fundamental components of these three phases as shown in Figure 5 below:
@@ III-C Deep Learning (DL)
In complex real-world problems, feature extraction can turn out to be extremely challenging using generic ML models. There may be hundreds of parameters that need to be learned and the outputs may not be linearly correlated to the inputs. So general ML models may not provide satisfactory performance in learning these problems. To facilitate mapping outputs to inputs, Neural Networks (NNs) [ai-nn] are widely used in ML frameworks. With the availability of a large amount of data, NNs have emerged as a key technology to be used in ML in the recent past. The learning process can be largely benefited from the introduction of NNs to deal with complicated large-scale problems. This learning process involving NNs to estimate the models is known as Deep Learning (DL) [ai-dl] which is a special important subset of ML.
NNs are inspired by the biological neural networks in the brain, more specifically the nervous system. To mimic the operation of the brain, the NNs are composed of multiple layers where each layer consists of multiple neurons followed by an activation function. Generally, the neurons in one layer are connected to the neurons in the adjacent layers. The connecting edges have weights that represent the relationship between the neurons. Each layer output can be viewed as some intermediate decisions which eventually result in the final output values. The weights are generally trained through a number of iterations using backpropagation algorithms [ai-bp]. Generally, the cost function associated with the model to calculate the difference between the predicted and actual outputs is not very simple, so we use different numerical methods like Gradient Descent [boydconvex], Stochastic Gradient Descent [sgd], Mini-Batch Stochastic Gradient Descent [mbsgd], Newton’s method [newton] etc. and so on to estimate the gradients of the cost function with respect to corresponding weights. At each iteration, the weights are updated by an amount based on these calculated gradients and a predefined learning rate. As we move towards the gradient descent direction, it helps us to reduce the cost at every iteration. In this manner, we can map the inputs to outputs through NNs.